OpenStudy (anonymous):

find the sum (2/3) + (1/8)

OpenStudy (anonymous):

0.79166666666

OpenStudy (anonymous):

\[\frac{2}{3}+\frac{1}{8}\]\[\frac{2 \times 8}{3 \times 8}+\frac{1 \times 3 }{8 \times 3}\]\[\frac{16}{24}+\frac{3}{24}\]\[\frac{19}{24}\]

OpenStudy (anonymous):

thanks